Bacteria X has a relative growth rate of 220 % (under ideal conditions). Some bacteria X are accidentally introduced into some egg salad. Three hours after contamination, there were 45,000 bacteria X in the egg salad.
Find the initial number of bacteria X introduced into the egg salad:

Answer by stanbon(75887)

Bacteria X has a relative growth rate of 220 % (under ideal conditions). Some bacteria X are accidentally introduced into some egg salad. Three hours after contamination, there were 45,000 bacteria X in the egg salad.
Find the initial number of bacteria X introduced into the egg salad:
--------------------------
Question: Do you mean 220% per hour?
A(t) = Ao(2.2)^t, where t is the number of hours.
-------------------
45000 = Ao(2.2)^3
Ao = 45000/2.2^3
Ao = 4226. 14
Rounding down:
Ao = 4226
====================
Cheers,
Stan H.